a table is 3 times as wide as it is long. the peremiter is 256.what is the width
Answer by checkley77(12844)

W=3L
2L+2W=256
2L+2*3L=256
2L+6L=256
8L=256
L=256/8
L=32 IS THE WIDTH.
W=3*32
W=96 IS THE LENGTH.
PROOF:
2*32+2*96=256
64+192=256
256=256